# Vision-Based CNN Prediction of Sunspot Numbers from SDO/HMI Images

### v1

@@ Abstract
Sunspot numbers constitute the longest and most widely used record of solar activity, with direct implications for space weather forecasting and heliophysical research. Traditional sunspot counting relies on visual inspection or algorithmic feature detection, both of which are limited by subjectivity, image quality, and methodological inconsistencies. Recent advances in deep learning, particularly convolutional neural networks (CNNs), enable the direct use of solar imagery for automated prediction tasks, reducing reliance on manual feature engineering. In this work, we present a supervised vision-based regression framework to estimate daily sunspot numbers from full-disk continuum images acquired by the Helioseismic and Magnetic Imager (HMI) onboard NASA’s Solar Dynamics Observatory (SDO). Images from 2011–2024 were paired with daily sunspot numbers from the SILSO Version 2.0 dataset of the Royal Observatory of Belgium. After preprocessing and augmentation, a CNN was trained to predict scalar sunspot counts directly from pixel data. The proposed model achieved strong predictive performance, with $R^{2}=0.986$ and RMSE = 6.25 on the test set, indicating close agreement with SILSO reference values. Comparative evaluation against prior studies shows that our approach performs competitively with, and in several cases outperforms, statistical and hybrid machine learning methods, while offering the novel advantage of bypassing explicit detection and manual feature extraction. Interpretability analyses using Grad-CAM and Integrated Gradients confirmed that the network consistently attends to sunspot regions when forming predictions. These results highlight the potential of deep vision-based approaches for operational solar monitoring, providing a scalable and automated pathway for real-time estimation of classical heliophysical indices. Future work should explore multimodal integration with magnetic field data and standardized benchmarking across solar cycles to strengthen operational applicability.
@@ 1 Introduction
Traditional manual methods for sunspot detection suffer from human subjectivity and inconsistencies, while automated techniques based on spatial filtering may compromise image resolution or distort physical sunspot properties (Sarsembayeva et al., 2021, 2025; Bornmann, 1990). Measurement accuracy has been shown to depend heavily on both the method and image quality, with systematic discrepancies noted between photographic records and sunspot drawing archives (Baranyi et al., 2001).
Technological advances have facilitated the development of high-resolution observational tools such as CCD imaging, spectroscopy, satellite platforms, and radio diagnostics, alongside tools for automated sunspot group area calculation, including software like S.C.A.T. (Solar Cycle Analyzer Tool) (Cedazo et al., 2020). Recent efforts have leveraged image processing and deep learning techniques to improve automatic sunspot detection, although accurate grouping of sunspots remains a challenge (Zhao et al., 2024). Modern computer vision methods, including Canny edge detection and topological analysis, have been applied to solar images from NASA’s Solar Dynamics Observatory (SDO) for feature identification (Sakpal, 2024; Rong et al., 2014; Cheng, 2012).
Recently, Acevedo (2024) proposed a machine learning method to predict sunspot numbers from solar imagery, leveraging topological feature extraction prior to modeling. In contrast, our work employs fully automated convolutional neural networks to directly map full-disk solar images to sunspot number predictions, removing the need for handcrafted feature design and enabling end-to-end regression learning. The rise of deep learning—especially Convolutional Neural Networks (CNNs)—has revolutionized solar physics by enabling the analysis of massive, high-dimensional datasets. SDO alone generates over 1.5 terabytes of data daily, rendering traditional processing approaches insufficient (Love et al., 2020). Deep learning models can extract hierarchical spatial features directly from raw observational data without the need for handcrafted input representations (Love et al., 2020; Chola and Benifa, 2022; Díaz Castillo et al., 2022).
@@
Advances in attention-based deep learning have also enabled full-disk solar flare prediction, overcoming the limitations of early models that focused only on central disk regions ($\pm 30^{\circ}$ to $\pm 45^{\circ}$) (Pandey et al., 2023; Huang et al., 2018; Li et al., 2020). These models have demonstrated average recall rates near 0.52 for near-limb flares beyond $\pm 70^{\circ}$ (Pandey et al., 2023).
@@
Computational constraints also arise in real-time applications, where deep models may be too resource-intensive for timely forecasts. Additionally, interpretability remains an issue: most CNNs behave as black boxes, limiting the ability of researchers to trust or explain model outputs. Finally, inconsistent evaluation metrics and benchmarking datasets across the literature hinder fair comparison between methods. Most applications to date focus on classification rather than continuous regression tasks, leaving a gap in the direct prediction of quantitative indices such as sunspot numbers or magnetic field strength.
@@
Our results show that deep learning techniques are not only capable of modeling visual solar patterns but also hold potential for quantitatively bridging observational solar imagery and fundamental heliophysical indices. This work thus contributes to expanding the toolkit for space weather forecasting by integrating automated vision-based approaches into the prediction of classical solar metrics.
@@ 2.1 Sunspot Number Data
Since 1981, the Solar Influences Data Analysis Center (SIDC) has served as the World Data Center for sunspot numbers, operating the SILSO (Sunspot Index and Long-term Solar Observations) project, which maintains the longest-running time series of solar activity measurements. The international sunspot number provided by SILSO is a vital reference for numerous scientific disciplines and supports over 500 institutional and individual users worldwide, including the International Astronomical Union and UNESCO.
@@
In this study, we used the daily total sunspot number covering the period from January 1st, 2011 to December 31st, 2024. This interval was selected to align with the availability of continuum solar images from the Helioseismic and Magnetic Imager (HMI) onboard NASA’s Solar Dynamics Observatory (SDO). The daily sunspot number was extracted as a single scalar value representing the global level of solar magnetic activity, which served as the target variable for our supervised regression task. In addition to its long-term stability and scientific acceptance, this dataset is particularly well suited as a ground truth reference because it encodes the aggregate visual manifestation of sunspots across the entire solar disk, making it directly comparable to the image-based features learned by convolutional neural networks.
@@
Figure 1: Daily sunspot number time series from SILSO (Version 2.0) between 2011 and 2024, including the smoothed solar cycle trend (https://www.sidc.be/SILSO/datafiles). This dataset served as the reference ground truth for model training and evaluation.
@@ 2.2 Solar Image Data from SDO/HMI
The Solar Dynamics Observatory (SDO), launched by NASA in 2010, continuously acquires full-disk images of the Sun across multiple spectral channels. For this study, we employed data from the Helioseismic and Magnetic Imager (HMI, https://soho.nascom.nasa.gov/data/REPROCESSING/Completed/), specifically the continuum intensity channel at 6173 Å. This wavelength samples the photospheric layer of the Sun and provides high-resolution images of the visible solar disk, where sunspots appear as dark regions due to their lower temperature relative to the surrounding photosphere. Continuum images are therefore particularly well suited for studies that require direct correspondence between sunspot morphology and the sunspot number index.
We selected one HMI continuum image per day over the 2011–2024 period, choosing the frame closest to 00:00 UTC to ensure daily consistency and temporal alignment with the SILSO sunspot index. Each raw image, originally recorded at $\times4096$ pixels, included metadata margins and peripheral artifacts. To standardize the dataset and prepare it for convolutional neural network (CNN) ingestion, we cropped the images to isolate the solar disk, removed extraneous borders, and downsampled them to $\times1024$ pixels using bilinear interpolation. Pixel intensities were normalized to the range [0,1] to stabilize the training process. In addition, data augmentation techniques such as random horizontal flips and small rotations (within $\pm 90^{\circ}$) were applied during training to enhance generalization and reduce overfitting. Each processed image was paired with the corresponding daily sunspot number from SILSO as its scalar regression target, thus enabling supervised learning.
@@ 2.3 Model Architecture and Training Procedure
We developed a supervised regression model based on a Convolutional Neural Network (CNN) architecture tailored for image-to-scalar prediction. After testing several configurations, we selected a model comprising six convolutional blocks with a combination of ReLU and Swish activations (Mercioni and Holban, 2020; Rahman et al., 2024), each followed by max pooling and batch normalization. The convolutional stack was followed by a global average pooling layer to reduce the spatial dimensions, and a fully connected dense layer with 128 neurons regularized with dropout and L2. The final output layer consisted of a single neuron with linear activation to produce a continuous-valued sunspot number prediction.
The model was trained using the Huber loss function (Gupta et al., 2020), which combines the robustness of the Mean Absolute Error (MAE) with the sensitivity of the Mean Squared Error (MSE), making it well-suited for handling outliers in solar activity. Optimization was performed using the Adam algorithm with a learning rate of $10^{-4}$ (Reddi et al., 2019). The dataset of 1890 image-label pairs was split chronologically into 80% for training and 20% for testing, preserving temporal consistency to prevent information leakage. Training was performed for a maximum of 100 epochs with early stopping based on validation loss, ensuring that the model converged efficiently while avoiding overfitting.
Model performance was evaluated using standard regression metrics, including MAE, Root Mean Squared Error (RMSE), and the coefficient of determination ($R^{2}$), calculated separately for the training and test sets. These metrics allowed us to assess both the absolute and relative accuracy of the model in capturing daily sunspot variability from solar images alone.
@@ 3 Results and Evaluation
The trained convolutional neural network demonstrated strong performance in estimating daily sunspot numbers from continuum solar images. Training curves showed rapid convergence within the first 15 epochs, followed by stabilization, with early stopping applied to prevent overfitting. Table 1 summarizes the predictive performance of the proposed CNN on both training and test splits. We report standard regression metrics—MAE, RMSE, and $R^{2}$—together with Pearson’s $r$, Spearman’s $\rho$, and bias, each accompanied by 95% bootstrap confidence intervals. The results indicate high agreement with the SILSO ground truth and low dispersion across the activity range, consistent with the patterns observed in Figures 36.
@@
Train 5.46 [5.02, 5.90] 7.58 [6.77, 8.59] 0.978 [0.972, 0.984] 0.990 [0.987, 0.992] 0.984 [0.978, 0.987] -0.823 [-1.471, -0.127]
Test 4.87 [4.20, 5.58] 6.25 [5.37, 7.17] 0.986 [0.980, 0.990] 0.993 [0.990, 0.995] 0.987 [0.978, 0.991] -0.406 [-1.539, 0.692]
The test split achieved an $R^{2}$ of $0.986$ with Pearson’s $r=0.993$ and Spearman’s $\rho=0.987$, indicating that the CNN captured both the magnitude and temporal variability of daily sunspot activity with high precision. Bias values remained close to zero in both splits, confirming the absence of systematic offsets between predictions and observations.
Figure 3 shows predicted versus observed sunspot numbers for both training and test sets. The regression lines closely follow the identity line, confirming that the model successfully tracks the underlying solar cycle variability, including ascending and descending phases. Minor dispersion appears in short-term fluctuations, particularly at higher values, but the overall temporal dynamics remain consistent.
@@
Residual and Bland–Altman plots (Figures 4 and 5) further illustrate the model’s behavior. Residuals are symmetrically distributed around zero with no evidence of heteroscedasticity, although a slight underestimation is observed during peak activity days. This reflects the limited number of extreme sunspot counts available in the training set.
@@
An error-stratified analysis by bins of solar activity (Figure 6) highlights that the prediction error increases with the mean sunspot number, as expected given the broader dynamic range near solar maxima. While MAE remains below $\approx 8$ even for the highest bins, this indicates that the relative difficulty of prediction grows during peak activity phases.
@@
Overall, these results confirm that vision-based regression is an effective strategy for estimating classical heliophysical indices directly from raw solar imagery, without explicit feature engineering or manual sunspot counting. The framework shows robustness across different levels of solar activity and provides a viable alternative for integrating automated deep learning pipelines into operational space weather monitoring.

### v2

@@ Abstract
Sunspot numbers constitute the longest and most widely used record of solar activity, with direct implications for heliophysical research and space-weather applications. Traditional sunspot counting relies on visual inspection and algorithmic feature-detection pipelines, which can be affected by observer-dependent choices, image quality, and methodological variability across implementations. Recent advances in deep learning, particularly convolutional neural networks (CNNs), enable the direct use of solar imagery for automated image-to-scalar regression, reducing the need for explicit, handcrafted feature design. In this work, we present a supervised vision-based framework to estimate the daily sunspot number from full-disk continuum images acquired by the Helioseismic and Magnetic Imager (HMI) onboard NASA’s Solar Dynamics Observatory (SDO). Images from 2011–2024 were paired with daily sunspot numbers from the SILSO Version 2.0 dataset maintained by the Royal Observatory of Belgium. After preprocessing and augmentation, a CNN was trained to infer a scalar sunspot number directly from pixel data at the observation time of each image. The proposed model achieved strong performance on an independent test split ($R^{2}$=0.964, RMSE=9.75, MAE=6.74), indicating close agreement with SILSO reference values across a broad activity range. In comparison with prior studies, we position this approach as a competitive and conceptually simple alternative for direct image-based estimation, complementing time-series forecasting models that target monthly means or smoothed indices. Interpretability analyses using Grad-CAM and Integrated Gradients indicate that the network consistently attributes relevance to sunspot-bearing regions when forming its estimates. These results highlight the potential of deep vision-based approaches for scalable solar monitoring and automated estimation of classical heliophysical indices. Future work should explore multimodal fusion with additional observables (e.g., magnetograms) and standardized cross-cycle benchmarks to strengthen robustness under changing solar conditions.
@@ 1 Introduction
Traditional manual methods for sunspot detection suffer from human subjectivity and inconsistencies, while automated techniques based on spatial filtering may compromise image resolution or distort physical sunspot properties (Sarsembayeva et al., 2021, 2025; Bornmann, 1990). Measurement accuracy has been shown to depend heavily on both the method and image quality, with systematic discrepancies noted between photographic records and sunspot drawing archives (Baranyi et al., 2001). Importantly, algorithmic detection pipelines also involve design choices (e.g., thresholds, morphology, and grouping rules) that can affect outputs; likewise, neural-network approaches depend on architectural and hyperparameter selections. In this work, the primary advantage of CNN-based regression is not the absence of methodological choices, but the ability to learn task-relevant representations directly from data in a consistent, end-to-end manner once the training protocol is fixed.
Technological advances have facilitated the development of high-resolution observational tools such as CCD imaging, spectroscopy, satellite platforms, and radio diagnostics, alongside tools for automated sunspot group area calculation, including software like S.C.A.T. (Solar Cycle Analyzer Tool; Cedazo et al., 2020). Recent efforts have leveraged image processing and deep learning techniques to improve automatic sunspot detection, although accurate grouping of sunspots remains a challenge (Zhao et al., 2024). Modern computer vision methods, including Canny edge detection and topological analysis, have been applied to solar images from NASA’s Solar Dynamics Observatory (SDO) for feature identification (Sakpal, 2024; Rong et al., 2014; Cheng, 2012).
Recently, Acevedo (2024) proposed a machine learning method to estimate sunspot numbers from solar imagery, leveraging topological feature extraction prior to modeling. In contrast, our work employs fully automated convolutional neural networks to directly map full-disk solar images to sunspot number estimates, removing the need for handcrafted feature design and enabling end-to-end regression learning. Note that we use the term “prediction” in the machine-learning sense of inferring a target value from input data at the same observation time (i.e., an image-to-index estimation), not as a forecast of future sunspot numbers. The rise of deep learning—especially Convolutional Neural Networks (CNNs)—has revolutionized solar physics by enabling the analysis of massive, high-dimensional datasets. SDO alone generates over 1.5 terabytes of data daily, motivating systematic, automated processing pipelines before the widespread adoption of deep learning (Love et al., 2020; Martens et al., 2012). Here, by “traditional” we refer to rule-based image processing and classical machine-learning pipelines that rely on manually engineered features; extensive benchmarking efforts for these methods exist in the literature (e.g., FLARECAST) (Georgoulis et al., 2021). Deep learning models can extract hierarchical spatial features directly from raw observational data without the need for handcrafted input representations (Love et al., 2020; Chola and Benifa, 2022; Díaz Castillo et al., 2022).
@@
Advances in attention-based deep learning have also enabled full-disk solar flare prediction, overcoming the limitations of early models that focused only on central disk regions ($\pm 30^{\circ}$ to $\pm 45^{\circ}$ in heliocentric angle from disk center, i.e., distance from disk center in angular coordinates) (Pandey et al., 2023; Huang et al., 2018; Li et al., 2020). These models have demonstrated average recall rates near 0.52 for near-limb flares beyond $\pm 70^{\circ}$ (Pandey et al., 2023).
@@
Computational constraints also arise in real-time applications, where deep models may be too resource-intensive for timely estimation. Additionally, interpretability remains an issue: most CNNs behave as black boxes, limiting the ability of researchers to trust or explain model outputs. Finally, inconsistent evaluation metrics and benchmarking datasets across the literature hinder fair comparison between methods. Most applications to date focus on classification rather than continuous regression tasks, leaving a gap in the direct prediction of quantitative indices such as sunspot numbers or magnetic field strength.
@@
Our results show that deep learning techniques are not only capable of modeling visual solar patterns but also hold potential for quantitatively bridging observational solar imagery and fundamental heliophysical indices. This work thus contributes to expanding the toolkit for space weather monitoring by integrating automated vision-based approaches into the estimation of classical solar metrics.
@@ 2.1 Sunspot Number Data
Since 1981, the Solar Influences Data Analysis Center (SIDC) has served as the World Data Center for sunspot numbers, operating the SILSO (Sunspot Index and Long-term Solar Observations) project, which maintains the longest-running time series of solar activity measurements. The international sunspot number provided by SILSO is a vital reference for numerous scientific disciplines and is widely used by the solar–terrestrial research community, including international institutions and organizations.
@@
In this study, we used the daily total sunspot number covering the period from January 1st, 2011 to February 29th, 2024. This interval was selected to align with the availability of continuum solar images from the Helioseismic and Magnetic Imager (HMI) onboard NASA’s Solar Dynamics Observatory (SDO). The daily sunspot number was extracted as a single scalar value representing the global level of solar magnetic activity, which served as the target variable for our supervised regression task. In addition to its long-term stability and scientific acceptance, this dataset is particularly well suited as a reference target for supervised learning because it encodes the aggregate visual manifestation of sunspots across the entire solar disk, making it directly comparable to the image-based features learned by convolutional neural networks.
@@
Figure 1: Daily sunspot number time series from SILSO (Version 2.0) between 2011 and 2024, including the smoothed solar cycle trend. This dataset served as the reference target for model training and evaluation.
@@ 2.2 Solar Image Data from SDO/HMI
The Solar Dynamics Observatory (SDO), launched by NASA in 2010, continuously acquires full-disk images of the Sun across multiple spectral channels.
For this study, we employed HMI continuum intensity observations at 6173 Å, accessed as publicly available full-disk quicklook JPEG images from the SOHO/SDO reprocessing archive (https://soho.nascom.nasa.gov/data/REPROCESSING/Completed/). This choice provides a uniform, readily reproducible image stream suitable for computer-vision ingestion, while preserving the photospheric contrast patterns associated with sunspot morphology. This wavelength samples the photospheric layer of the Sun and provides high-resolution images of the visible solar disk, where sunspots appear as dark regions due to their lower temperature relative to the surrounding photosphere. Continuum images are therefore particularly well suited for studies that require direct correspondence between sunspot morphology and the sunspot number index.
We selected one HMI continuum image per day from 2011-01-01 to 2024-02-29, choosing the frame closest to 00:00 UTC to ensure daily consistency and temporal alignment with the SILSO sunspot index. Each raw image, originally recorded at $\times4096$ pixels, included metadata margins and peripheral artifacts. To standardize the dataset and prepare it for convolutional neural network (CNN) ingestion, we cropped the images to isolate the solar disk, removed extraneous borders, and downsampled them to $\times512$ pixels using bilinear interpolation. This preprocessing enforces a fixed input geometry for the CNN while retaining the full-disk context. Pixel intensities were normalized to the range [0,1] to stabilize the training process. In addition, data augmentation techniques such as random horizontal flips, small rotations (within $\pm 10^{\circ}$), and isotropic rescaling within $\pm 3\%$ were applied during training to enhance generalization and reduce overfitting. Each processed image was paired with the corresponding daily sunspot number from SILSO as its scalar regression target, thus enabling supervised learning. The resulting dataset contains 4778 daily images, split into 3822 for training and 956 for testing.
@@ 2.3 Model Architecture and Training Procedure
We developed a supervised regression model based on a Convolutional Neural Network (CNN) architecture tailored for image-to-scalar estimation. The final model consists of a hierarchy of convolutional stages using $3\times 3$ kernels and Swish activations (Mercioni and Holban, 2020; Rahman et al., 2024). Each stage applies two convolutional layers followed by batch normalization and spatial downsampling via max pooling, progressively increasing the number of feature channels (32, 64, 128, 256, and 512) while reducing spatial resolution. The convolutional backbone is followed by a global average pooling layer and a lightweight fully connected head (Dense 256 with dropout, Dense 128), ending in a single linear neuron to output a continuous sunspot number estimate.
The model was trained using the Huber loss function (Gupta et al., 2020), which combines the robustness of the Mean Absolute Error (MAE) with the sensitivity of the Mean Squared Error (MSE), making it well-suited for handling outliers in solar activity. Optimization was performed using the Adam algorithm with a learning rate of $10^{-4}$ (Reddi et al., 2019). The dataset comprises 4778 daily image–label pairs and was split chronologically into training (3822 images) and testing (956 images) subsets to preserve temporal consistency and prevent information leakage. Model selection and early stopping were driven by a validation subset carved out from the training period (using a fixed hold-out fraction of the training data), and training was run for up to 100 epochs with early stopping on validation loss to mitigate overfitting.
Model performance was evaluated using standard regression metrics, including MAE, Root Mean Squared Error (RMSE), and the coefficient of determination ($R^{2}$), calculated separately for the training and test sets. These metrics allowed us to assess both the absolute and relative accuracy of the model in capturing daily sunspot variability from solar images alone. Training was performed on a consumer-grade NVIDIA GPU (GeForce MX250, 2 GB VRAM) using TensorFlow/Keras, with typical training runs completing within a few hours depending on early stopping.
Figure 3 summarizes the final network topology. The backbone follows a VGG-like pattern of stacked $3\times 3$ convolutions with batch normalization and periodic max-pooling, which progressively reduces the spatial resolution (from $512\times 512$ down to $16\times 16$) while increasing the number of feature channels (32 to 512). This design encourages the model to learn low-level contrast and edge cues in early layers and increasingly abstract, multi-scale representations of active-region morphology at deeper stages. A Global Average Pooling (GAP) layer then aggregates the final feature maps into a compact 512-dimensional descriptor, which is mapped to a scalar estimate through a lightweight fully connected head (Dense 256 with dropout, Dense 128, and a final linear neuron). This combination retains full-disk context, limits the number of parameters in the regressor head, and supports interpretability analyses based on gradient attribution.
Figure 3: Schematic overview of the final CNN regression architecture. The model processes $512\times 512$ three-channel HMI continuum quicklook images through stacked $3\times 3$ convolutional layers with batch normalization and max pooling, followed by Global Average Pooling (512) and a dense regression head (Dense 256 + dropout, Dense 128, Output 1).
@@ 3 Results and Evaluation
3.1 Quantitative results, regression performance and residual analysis
The trained convolutional neural network demonstrated strong performance in estimating daily sunspot numbers from continuum solar images. Training curves showed rapid convergence within the first 15 epochs, followed by stabilization, with early stopping applied to prevent overfitting. Table 1 summarizes the predictive performance of the proposed CNN on both training and test splits. We report standard regression metrics—MAE, RMSE, and $R^{2}$—together with Pearson’s $r$, Spearman’s $\rho$, and bias, each accompanied by 95% bootstrap confidence intervals. The results indicate high agreement with the SILSO ground truth, with an error structure that remains stable overall and increases moderately toward high-activity regimes, consistent with Figures 47.
@@ 3.1 Quantitative results, regression performance and residual analysis
Train 6.82 [6.61, 7.03] 9.64 [9.38, 9.90] 0.964 [0.961, 0.966] 0.983 [0.982, 0.984] 0.984 [0.982, 0.984] 2.11 [1.84, 2.41]
Test 6.74 [6.29, 7.19] 9.75 [9.18, 10.35] 0.964 [0.960, 0.969] 0.983 [0.981, 0.985] 0.982 [0.979, 0.984] 1.77 [1.15, 2.36]
Both splits achieved comparable performance (Table 1), with overlapping confidence intervals across all metrics. On the test split, the model reached $R^{2}=0.964$ with Pearson’s $r=0.983$ and Spearman’s $\rho=0.982$, indicating strong agreement with the SILSO reference. The mean bias is slightly positive in both splits, suggesting a small overall tendency to overestimate sunspot numbers on average, while remaining within narrow confidence bounds.
Figure 4 shows predicted versus observed sunspot numbers for both training and test sets. The regression lines closely follow the identity line, indicating accurate calibration across most of the activity range. Although the model processes each daily image independently (i.e., it is not a temporal model), the time-ordered sequence of daily estimates reproduces the observed day-to-day variations of solar activity. Dispersion increases at higher sunspot numbers, consistent with the larger intrinsic variability during active periods.
@@
Residual and Bland–Altman plots (Figures 5 and 6) further illustrate the model’s error structure. Residuals remain centered close to zero, but their spread increases with activity level, indicating mild heteroscedasticity. The Bland–Altman analysis shows a small positive mean difference (consistent with the overall bias), with wider limits of agreement at higher sunspot numbers. In the most active regimes, a slight regression-to-the-mean behavior is observed, with occasional underestimation of extreme values, likely reflecting the reduced availability of high-activity examples.
@@
An error-stratified analysis by bins of solar activity (Figure 7) shows that prediction error increases with the mean sunspot number. On the test split, MAE rises from $\approx 2$ at very low activity to $\approx 11$ in the highest-activity bin, consistent with both the broader dynamic range near solar maxima and the relative scarcity of extreme-activity days. Notably, the sign of the bias can change at the upper end, indicating a tendency to slightly underpredict the most extreme values while mildly overpredicting intermediate activity levels.
@@
Taken together, these diagnostics indicate that the proposed CNN provides a well-calibrated mapping from full-disk continuum imagery to the daily sunspot number, with nearly identical behavior on training and test splits. The remaining discrepancies are largely structured rather than random: error variance increases with activity level and the most extreme cases exhibit a mild regression-to-the-mean effect, which is consistent with both the wider dynamical range near solar maxima and the smaller number of high-activity examples available for learning. Importantly, because the model operates on each daily image independently, the apparent time evolution in the outputs reflects the chronological ordering of independent estimates rather than an explicit temporal component in the architecture. These patterns motivate the interpretability analyses in the next section, where we examine which solar-disk regions and visual structures drive the model’s estimates across different activity regimes. Overall, the results support vision-based regression as a practical pathway to estimate classical heliophysical indices from full-disk imagery in a fully automated manner.